5 The following financial statement information is from five separate companies: Company Company Company Company Company Part 5 of 5 December 31, 2008 $38,000 $29,640 $24,320 $67640 $103,740 Assets Liabilities 31,160 20,74813,132 46,671 2 December 31, 2009 74,620 113160 ? 20,073 13.460 35,817 89,396 41,000 29,520 Assets Liabilities During year 2009 eBook 6,500 7,516 0 11,000 6,0001,4009,750 Owner investments Net income (loss) Owner cash withdrawals 1,053 10,834 9,110 3,5002.000 Print 5,875 Requirement 5: Answer the ollowing questions about Company E (Omit the $ sign in your response): a. What is the amount of equity on December 31, 2008? b. What is the amount of liabilities on December 31 2008? Mc Graw Hill < Plex 5 of 15 İİİ Next > O Type here to search
Answer #1

Company E :

Ending owner's equity = 113160-89396 = 23764

a) Beginning owner's equity (2008) = 23764-6500-7516+11000 = 20748

b) Beginning liabilities (2008) = 103740-20748 = $82992